A television loop antenna has a diameter of 2.1m. The magnetic field of a TV signal is normal to the plane loop and, its magnitude is 0.42 T. Calculate the magnetic flux produced.
1
Expert's answer
2019-10-28T11:35:44-0400

We determine the area of the antenna

"S=\pi\cdot (D^2)/4 = 3.14\cdot (2.1^2)/4=7.544 m^2"

Then the magnetic flux is

"\Phi = В\cdot S = 0.42\cdot 7.544 =3.168 Wb"